### v1

# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

Abstract
Batch normalization (BN) is a ubiquitous operation in deep neural networks used primarily to achieve stability and regularization during network training. BN involves feature map centering and scaling using sample means and variances, respectively. Since these statistics are being estimated across the feature maps within a batch, this problem is ideally suited for the application of Stein’s shrinkage estimation, which leads to a better, in the mean-squared-error sense, estimate of the mean and variance of the batch. In this paper, we prove that the Stein shrinkage estimator for the mean and variance dominates over the sample mean and variance estimators in the presence of adversarial attacks when modeling these attacks using sub-Gaussian distributions. This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and use it in image classification (segmentation) tasks with and without adversarial attacks. We present SOTA performance results using this Stein corrected batch norm in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
1 Introduction
Deep Neural Networks are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and sciences. In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of deep neural networks. Deep network training is well known to be time intensive and unstable. Batch normalization (BN) is a particularly useful tool in achieving numerical stability and regularization during network training. BN typically involves centering and scaling the feature maps. The centering and scaling operations involve computing the sample means and variances, respectively, from the samples/members in the batch. This task is ideally suited for the application of Stein’s shrinkage method, which is well known in the field of statistics. Before delving into the Stein’s shrinkage method, let us briefly review the widely popular batch normalization in deep learning.
Batch normalization was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift Ioffe-ICML15 . Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) ba2016layernormalization and Transformers, spectral normalization in generative adversarial networks (GANs) Miyato-ICLR18 , and many others. Improvements over BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training has also been proposed in Zhang_2021_CVPR and batch normalization preconditioning in Lange-JMLR2022 , wherein the pre-conditioner is constructed using batch statistics. These normalization algorithms do not address the effectiveness of improvements in stability or convergence of training in the presence of adversarial attacks.
In this paper, we address the problem of developing a BN that leads to better prediction accuracy from the deep network used in classification in the presence of adversarial attacks. Adversarial attacks have been modeled via additive bounded perturbations MauryaArxiv22 . Some of the popular adversarial attack methods that are popular and satisfy the bounded perturbation condition are FGSM (fast gradient sign method) szegedy2014 and an iterative variaint of PGD (projected gradient descent). Numerous articles exist in the literature that depict ways to overcome a drop in the accuracy of classifiers in the presence of adversarial attacks. Our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, however, to adapt an existing shrinkage estimator specifically, the well known James-Stein shrinkage estimator and prove that it retains its dominance over the sample mean and variance in the presence of adversarial attacks. Dominance here refers to smaller mean squared error (MSE) – between the estimate and the true value – when compared to the sample mean and variance. This smaller MSE implies a better BN in each layer and as we demonstrate empirically, a better stability and prediction accuracy from the deep network. By making BN stable in the presence of adversarial attacks this work could have potential applications that might have societal impact.
1.1 A Brief on the James-Stein Shrinkage Estimator
In the following we present a brief introduction to the James-Stein (JS) shrinkage estimator and motivate it’s aptness in the context of batch normalization in deep neural networks. This will provide the necessary background and motivation for our work presented in this paper.
Shrinkage estimation is a parameter estimation technique that is primarily used for variance reduction and to improve parameter estimation accuracy by shrinking the parameter towards a fixed target. For instance, shrinking the estimate of the mean of a Gaussian toward zero. It is especially useful in high-dimensional and scarce data settings. The shrinkage, of course, leads to a biased estimate, but with smaller variance, which is sought after in many applications. The James-Stein (JS) shrinkage estimator was proposed in the field of statistics as an estimator of the mean for samples drawn from a Gaussian distribution and has been shown to dominate the maximum likelihood estimator (MLE) in terms of the risk (mean-squared error) James-Stein61 . Dominance here refers to the risk associated with the JS-estimator being at most as much as that of the MLE and that there exists at least one set of parameter values for which it is strictly less. More precisely, if $\mathbf{X}$ is a $p$=dimensional random vector whose mean is to be estimated and it is assumed that $\mathbf{X}\sim N(\mathbf{\mu},\mathbf{I})$, then, $E(||\mathbf{\mu}^{JS}-\mathbf{\mu}||^{2})\leq E(||\mathbf{\mu}^{MLE}-\mathbf{\mu}||^{2})$. It is important to note that the JS-estimator improves the overall risk of estimating the mean (parameter) vector and not the risk of the individual elements of the mean vector.
At the foundation of the JS-shrinkage estimator lies what is known as Stein’s (identity) lemma, which characterizes the normal distribution. Stein’s (identity) lemma is also at the foundation of the Stein’s method which is used for distributional comparisons. We will not discuss this latter well-known method in statistics here and instead refer the interested reader to a recent comprehensive survey Anastasiou23 .
Since its inception, the JS-shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family Brown66 , exponential family Hudson78AOS , Gamma distribution DasGupta86AOS , distributions with quadratic varaince function XianchaoXie-Brown-AOS16 and FathiAOS2022 for relaxing the Gaussian assumption in high dimensions. For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to Kubokawa24 . In addition, more recently, Stein’s shrinkage estimate has been generalized to distributions on manifold-valued random variables Yang-Doss-Vemuri22 .
1.2 Relevance of JS-estimator to BN
Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to batch normalization in deep neural networks and how it can be applied to improve the stability and accuracy of predictions. A wide range of normalization techniques have been developed to facilitate the training of deep neural networks by normalizing intermediate feature distributions and improving the stochastic optimization dynamics. One of the earliest and most influential approaches is BN, introduced in Ioffe-ICML15 , which normalizes features using statistics computed over the entire mini-batch. This method was originally motivated by the desire to reduce internal covariate shift, though subsequent work has highlighted its strong impact on the optimization landscape conf/nips/SanturkarTIM18 as a key factor behind its effectiveness. In the following, we give a brief and precise description of different types of normalization that are in vogue in literature.
Let $X\in\mathbb{R}^{N\times C\times H\times W}$ denote the feature tensor at a given layer of a deep neural network, where $N$ – batch size, $C$ – number of channels (feature maps), $H,W$ – spatial dimension. Formally, many normalization techniques follow a common template. Each element $x_{i}\in X$ is normalized as: $\hat{x_{i}}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}}$, where $i=(i_{N},i_{C},i_{H},i_{W})$ is a multi-index ranging over the tensor dimensions and $\mu_{i},\sigma^{2}_{i}$ are the empirical mean and variance computed over a subset $S_{i}\subseteq X:$ $\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\ \ \sigma_{i}^{2}=\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2}.$ The choice of $S_{i}$ determines the normalization scheme and the available choices are the following: (i) Batch Normalization (BN)  Ioffe-ICML15 : $S_{i}^{BN}={(n^{\prime},h^{\prime},w^{\prime})|n^{\prime}\in[1,N],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed over all samples in the batch and all spatial positions, but separately for each channel $i_{c}.$ (ii) Layer Normalization (LN)  ba2016layernormalization : $S_{i}^{LN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in[1,C],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed independently for each sample (fixed $i_{N}$) across all channels and spatial locations. (iii) Instance Normalization (LN)  ulyanov2017instancenormalizationmissingingredient : $S_{i}^{IN}={(h^{\prime},w^{\prime})|h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Normalization is per-sample and per-channel — only over the spatial dimensions. (iv) Group Normalization (GN)  wu2018groupnormalization : $S_{i}^{GN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in C_{\text{group}},h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Channels are split into groups; statistics are computed over spatial dimensions and within the channel group. After normalization, a learnable affine transformation is applied to preserve representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
While normalization methods like LayerNorm, InstanceNorm, and GroupNorm have been introduced to mitigate batch-size dependence, BatchNorm remains a central component in many modern architectures and continues to attract interest both in theory and practice. A key observation is that BatchNorm estimates a large number of mean and variance values simultaneously — one pair for each channel. This can be naturally viewed as a multitask estimation problem: we are estimating a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$ each based on a finite number of observations.
In such settings, James–Stein shrinkage estimators offer a statistically grounded way to improve estimation by borrowing strength across tasks — in this case, across channels. The classical result suggests that when estimating multiple parameters jointly under squared loss, shrinking individual estimates toward their grand mean can reduce overall error. Applying this idea in the context of BatchNorm could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
Having established adequate motivation for application of Stein’s shrinkage to BN in a deep network, we are now ready to narrate the rest of the story. But before that, a word on what is to follow. Section 2 delves into the presentation of the appropriate shrinkage formulas for the BN. This is followed by the presentation of theoretical results on the dominance of Stein’s shrinkage formulas in the presence of adversarial attacks in Section 3. In Section 4, we present experimental results that compare the performance of the proposed shrinkage to SOTA on several benchmark data sets. Finally in Section 5, we present the discussion and conclusions.
2 Stein’s Shrinkage Applied to Batch Normalization
In this section, we apply the Stein’s shrinkage to the BN and present the formulas for calculation of the batch normalization in a deep neural network. We will limit ourselves primarily to convolutional neural networks (CNNs) but the formulas can be applied to any architecture with some minor changes.
Let $X\sim\mathcal{N}_{p}(\theta,I_{p})$, where $\theta\in\mathbb{R}^{p}$ is an unknown parameter vector and $I_{p}$ is the $p$-dimensional identity matrix. The James Stein shrinkage estimator of the unknown parameter $\theta$ is given by, $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X$. It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator but is obviously a biased estimator. However, a smaller mean squared error is desirable in many applications. Most recently, JS-estimator for the calculation of the mean in BN was used in khoshsirat2023improvingnormalizationjamessteinestimator . However, the authors in khoshsirat2023improvingnormalizationjamessteinestimator used the same JS-estimator formula to shrink the mean as well as the variance in BN. Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the sample variance is not normally distributed but instead follows a scaled chi-squared distribution - a special case of the Gamma distribution. To address this, we adopt a principled alternative.
More specifically, for mean correction, we adopt the James-Stein shrinkage under the assumption that individual elements of the feature tensor in the BN are normally distributed and homoscedastic. The homoscedasicity assumption simplifies our calculations and we will address relaxing this assumption in future. Under this assumption, the sample mean is normally distributed, and the sample variance follows a chi-squared distribution (i.e., a Gamma distribution): $\hat{\sigma}_{i}^{2}\sim\frac{\sigma_{i}^{2}}{n}\cdot\chi^{2}_{n-1}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$, which enables us to apply shrinkage to the scale parameter of the Gamma distribution DasGupta86AOS . Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ independently for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$. Then JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right]$ (1)
To incorporate the James-Stein correction into Batch Normalization, we compute the statistics for each channel and apply the shrinkage formulas as outlined below. Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions given by:
$\mu_{C}=\frac{1}{NHW}\sum_{n,h,w}x_{n,c,h,w},\quad\sigma^{2}_{C}=\frac{1}{NHW}\sum_{n,h,w}(x_{n,c,h,w}-\mu_{C})^{2}.$
We then apply shrinkage to these estimates. The James-Stein corrected mean and variance are:
Stein Shrinkage Forumals for Mean & Variance $\mu^{\mathrm{JS}}_{C}$ $=\left(1-\frac{(c-2)\sigma^{2}_{\mu_{C}}}{\|\mu_{C}\|^{2}_{2}}\right)\mu_{C},\quad{\sigma_{C}^{2}}^{\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+\tilde{c}\cdot V,$ (2) $\text{where }V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad\tilde{c}\in\left[0,\dfrac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$
Here, $\sigma^{2}_{\mu_{C}}$ denotes the estimated variance of the corresponding vector $\mu_{C}$.
Finally, the normalized output is given by:
Final Normalized (scaled and shifted) Output $\hat{x}=\frac{x-\mu^{\mathrm{JS}}}{\sqrt{{\sigma^{2}}^{\mathrm{JS}}+\epsilon}},\quad y=\gamma\hat{x}+\beta.$ (3)
Equations 2 and 3 respectively are the formulas we advocate for use in BN corrections and will be used in all our experiments. We are now ready to present the theoretical results of dominance of these shrinkage estimators in the presence of adverserial attacks.
3 Dominance of Stein Shrinkage in the Presence of Adverserial Attacks
In this section, we present the main theoretical contribution of our work: the James–Stein estimator continues to exhibit dominance even under adversarial perturbations. We begin by formalizing how adversarial attacks can be modeled probabilistically. Then, we introduce the necessary preliminaries and outline the main ideas behind the proof.
We model adversarial perturbations using a sub-Gaussian distribution, which is well-motivated both practically and theoretically. Adversarial perturbations are typically bounded in magnitude (e.g., under an $l_{p}$-norm constraint) to ensure the model remains learnable — a standard assumption in the literature cohen2019certifiedadversarialrobustnessrandomized ; zhai2019adversariallyrobustgeneralizationjust . Since bounded random variables are inherently sub-Gaussian, this provides a natural modeling choice. Additionally, the sub-Gaussian model enables a tractable analysis via the use of concentration inequalities and allows one to work with a well-defined variance proxy. Following prior workMauryaArxiv22 , we assume that each adversarial perturbation vector has coordinates bounded in $[-\varepsilon,\varepsilon]$ and can be represented as a zero-mean sub-Gaussian vector with variance proxy $2\varepsilon^{2}$: $Y\sim\text{subG}(2\varepsilon^{2}),\quad\text{i.e., for all }v\in\mathbb{R}^{p}\text{ with }\|v\|=1,\ \ \text{and}\ \ \mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\frac{2\varepsilon^{2}\lambda^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$ This assumption allows us to model the observed vector as the sum of noise-free data and adversarial noise: $Z=X+Y,\ \ \ X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$ For simplicity we assume that $X$ and $Y$ are independent, which is currently a limitation of our model. Other noise models that are data dependent will be considered in our future work.
3.1 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Normally Distributed Random Variables
Consider the problem of estimating a parameter vector $\theta\in\mathbb{R}^{p}$ based on corrupted observations $Z=X+Y,$ where $X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I_{p})$ and $Y\in\mathbb{R}^{p}$ is independent, mean-zero sub-Gaussian vector with a known proxy variance $2\varepsilon^{2}.$ Our goal is to compare the risks (expectation of the loss) of the naive estimator $\hat{\theta}^{\mathrm{MLE}}$ and the JS-estimator
$\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{(p-2)\hat{\sigma}^{2}}{\|Z\|^{2}_{2}}\right)Z.$
under squared error loss. It is well-known that $\hat{\theta}^{\mathrm{JS}}$ strictly dominates $\hat{\theta}^{\mathrm{MLE}}$ when $Y=0$ and $p\geq 3$ James-Stein61 . We aim to show that this dominance remains valid even under additive sub-Gaussian noise, provided the proxy variance $2\varepsilon^{2}$ is sufficiently small.
We first observe that it suffices to consider the case $\sigma^{2}=1$ The result for general $\sigma^{2}$ then follows by a simple scaling argument. Our first key result is the following dominance theorem for the Gaussian random variable:
JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gaussian Model
Theorem 3.1.
Let $\theta\in\mathbb{R}^{p}$ be an unknown parameter vector. Suppose we observe $Z=X+Y$, where $X\sim\mathcal{N}_{p}(\theta,I_{p})$ and $Y\in\mathbb{R}^{p}$ is an independent, mean-zero sub-Gaussian vector with parameter $2\varepsilon^{2}$, i.e., for all unit vectors $v\in\mathbb{R}^{p}$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\varepsilon^{2}\lambda^{2}\right)$. Define the JS-estimator by $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|Z\|^{2}}\right)Z,$ and the maximum likelihood estimator by $\hat{\theta}^{\mathrm{MLE}}=Z$. Then, for all $\theta\in\mathbb{R}^{p}$ and for all $p\geq 3$ the JS-estimator strictly dominates the MLE in mean squared risk: $R(\hat{\theta}^{\mathrm{JS}},\theta)<R(\hat{\theta}^{\mathrm{MLE}},\theta).$
We now provide a sketch of the proof and refer the reader to the Appendix for more details.
Sketch of Proof.
We can rewrite the risk of the JS-estimator as:
$R(\hat{\theta}^{\mathrm{JS}},\theta)=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right).$ (4)
The last two terms in equation 4 can be rewritten as:
$-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\right)$
Further, by applying Cauchy–Schwarz and Jensen’s inequalities we prove (in the Appendix) that: $\mathbb{E}\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$ Hence, the risk of the James–Stein estimator is strictly less than that of the MLE for $p\geq 3$ completing the proof. ∎
We are now ready to address the issue of shrinkage estimation of the variance parameter in the above normally distributed random variable with additive sub-Gaussian perturbation.
3.2 Dominance of the JS-estimator under Sub-Gaussian Perturbations for Gamma Distributed Random Variables
In many applications, empirical variances follow a scaled chi-squared distribution, which is a special case of the gamma distribution. Given observations $x_{ij}$, define the empirical variance: $\hat{\sigma}_{i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(x_{ij}-\bar{x}_{i})^{2}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$. Thus, estimating $\sigma_{i}^{2}$ reduces to estimating the scale parameter $\beta_{i}$ in a gamma distribution with fixed shape parameter $\alpha$. For this task, we can use the following theorem from literature:
Theorem 3.2 (JS-Estimator Shrinkage Dominance for the Gamma Model DasGupta86AOS ).
[…]
Now, applying the above theorem to empirical variances, we obtain: $\hat{\sigma}_{i}^{2,\mathrm{JS}}=\frac{n}{n+1}\hat{\sigma}_{i}^{2}+cV,\quad V=\left(\prod_{j=1}^{p}\hat{\sigma}_{j}^{2}\right)^{1/p},\quad c\in\left[0,\frac{4n(p-1)}{(n+1)((n-1)p+2)}\right].$ We now present our second key result namely, that this dominance persists under additive sub-Gaussian noise for the Gamma distributed random variable.
JS-Estimator Dominance under Sub-Gaussian Perturbations for the Gamma Model
Theorem 3.3.
[…]
Remark 3.4.
The admissible range for $c$ under sub-Gaussian perturbations includes the classical interval, so standard shrinkage remains valid even in noisy settings.
Sketch of Proof.
[…]
We begin by expressing the risk of $\hat{\beta}^{\,\mathrm{JS}}$ as:
[…]
Thus, the difference in risks depends on the sign of the second term in the above expression expressed as: $c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right).$ To show that this quantity is negative, we analyze the sign depending on whether $c>0$ or $c<0$, and whether the linear term dominates the quadratic term. We consider two main cases:
  • Case 1:
    $c<0$ and
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (5)
  • Case 2:
    $c>0$ and
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (6)
In both cases outlined above, the proof fundamentally relies on an application of Stein’s lemma for the Gamma distribution given below. Since each $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ with known shape and rate parameters, we can invoke the Stein’s identity given in the following lemma:
Lemma 3.5 (Stein’s Lemma for the Gamma Distribution).
Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty$. Then the following (Stein’s) identity holds: $\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)]$.
We apply Stein’s identity to the function $h(x)=V$, where $V$ is the geometric mean defined by $V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}$. Using the Stein’s identity for the Gamma distributed random variable, which is given by, $\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$ we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ and get: $\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\left(\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}\right).$ Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get: $\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
Depending on the sign and structure of the risk difference (as in the two main cases above), we proceed to estimate the relevant expectations using various inequalities — such as Bernstein’s inequality (to control sub-exponential terms like $W_{i}$), Jensen’s inequality, Cauchy–Schwarz, and the arithmetic–geometric mean inequality. These tools allow us to bound the expressions for a given choice of $c$ and derive the dominance for the JS-estimator of the variance parameter of the Normally distributed random variable under sub-Gaussian perturbations.
Further technical details and precise bounds are deferred to the appendix. ∎
4 Experiments
In this section, we present several experimental results demonstrating the performance of BN with our proposed shrinkage estimators in both image classification and segmentation tasks, under standard and adversarial settings. To simulate adversarial attacks, we added noise samples drawn from a sub-Gaussian distribution with a fixed variance. The sub-Gaussian density is given by $f(x)=\frac{1}{2\sqrt{2}\pi\sigma}\left(\frac{x^{2}}{2\sigma^{2}}+\frac{1}{4}\right)^{-1},$ derived from a mixture of Lévy and Gaussian distributions, and the samples were drawn using the inverse transform sampling technique. GEORGIOU20063061 . For classification, we employed the ResNet-9 he2015deepresiduallearningimage architecture on the CIFAR-10  cifar10 dataset, as well as a 3D CNN model on the PPMI dataset ppmi-data . For object segmentation, we used the HRNetV2 wang2020deep model applied to the Cityscapes datasetCordts2015Cvprw . Since BN is known to become unstable when the batch size is small, we report results for varying batch sizes and sub-Gaussian noise levels. Our goal is to demonstrate that our shrinkage-based method is more robust to adversarial perturbations and offers greater overall stability. We remind the reader that our goal here is not to improve the SOTA results on these data in the literature, but to demonstrate the performance gains from the use of BN in the presence (absence) of adversarial attacks with the correct use of Stein shrinkage compared to vanilla BN and other forms of shrinkage applied to BN. Throughout all the experiments, we avoided training-time tricks and focused on obtaining a honest comparison.
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
Standard (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
Standard (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
Our (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat khoshsirat2023improvingnormalizationjamessteinestimator (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
Lasso (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
Lasso (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
Lasso corrected (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
Ridge (32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
Ridge (64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
Ridge (128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
Table 1: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
In the experiments, we compare our approach mainly with that of Khoshsirat et al. khoshsirat2023improvingnormalizationjamessteinestimator , aiming to show that theoretically correct estimation of mean as well as variance shrinkage significantly improves performance, particularly in small batch regimes. In addition, with the correct variance shrinkage, one can achieve higher accuracy in the presence of adversarial perturbations. We also tested a variant using only the mean shrinkage estimator and no shrinkage on variance. Additionally, we included models regularized with Lasso and Ridge methods for comparison, as these are commonly used shrinkage methods. Detailed formulas for these estimators and complete experimental setup are provided in the Appendix B. Note that in all the tables depicting the results, we have highlighted the best performing BN method (ours) for small batch sizes that are preferred from a computational efficiency perspective as well as when data are scarce.
4.1 Application to Cifar10 Data
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. The results in Table 1 confirm that our proposed BN significantly outperforms the baselines in noisy settings across all batch sizes. Although standard BN variants quickly degrade under noise, our method maintains substantially higher accuracy, especially with moderate to large noise levels $(10-30\%).$ Compared to the method in khoshsirat2023improvingnormalizationjamessteinestimator , our approach shows consistent improvement due to the correct development and use of variance shrinkage. The method with only mean correction (see Appendix) performs worse than our full model, highlighting the importance of variance correction. Finally, Lasso and Ridge-based approaches do not improve noise performance, showing behavior similar to the standard BN.
4.2 Application to Cityscapes Data
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
Standard (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
Our (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
Lasso (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
Lasso (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
Ridge (8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
Ridge (16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset—a more challenging task involving high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. As in the previous experiment, we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. This setup is particularly relevant for segmentation, where maintaining accurate local feature distributions is crucial for pixel-level predictions. As shown in Table 2, standard BN fails to retain meaningful performance even under mild noise (e.g., below $10$ mIoU score at $3\%$ corruption). In contrast, our proposed method and its mean corrected variant (see the appendix) significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings. Regularization-based approaches (Lasso and Ridge) show negligible improvement under noise and have a behavior similar to that of standard BN.
4.3 Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN method on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline (see Appendix B) using the FSL software FSLpack . DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local Gaussian fit. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization FORTIN2017149 is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used. The data set is highly imbalanced and contains FA maps from 360 Control and 130 PD brain scans. This imbalance is addressed using mixup-basedDBLP:journals/corr/abs-1710-09412 augmentation with sampling bias favoring the smaller sized class (PD). Using the same setup as in previous experiments and a 5-fold cross-validation to compare 3D CNN with different variants of 3D BN, we get results summarized in the table 3. These results further support our earlier observations, even in this challenging classification task.
BN Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge (10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge (16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on PPMI data. Values depict mean accuracies $\pm$ standard deviation.
5 Conclusions
In this paper, we presented a novel theory involving the proof of dominance of the Stein’s shrinkage estimators of mean and variance of a Gaussian random variable in the presence of adversarial attacks. We applied Stein’s shirnkage to BN in a variety of deep networks and demonstrated superior performance, in comparison to other BN methods, on three benchmark data sets. One possible explanation for the improved performance is the connection between model performance under attack and the Lipschitz properties of the network DBLP:journals/corr/abs-2003-02460 . Since BN is known to affect the Lipschitz constant conf/nips/SanturkarTIM18 , it would be interesting to investigate the influence of our BN method on this aspect in future work.
6 Acknowledgments
This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099. (Corresponding author: Baba C. Vemuri.)
Appendix A Technical Appendices and Supplementary Material
In this Appendix, we present the detailed proofs of the two key theorems 3.1 and 3.3 and additional experiments showing Stein’s correction applied only to the mean and not the variance. These experimental results are depicted here in this Appendix as additional rows added in to the three tables appearing in the main body of the paper. Also included are further experimental details.
Theorem 3.1 Proof:.
We denote MLE estimator as $\hat{\theta}^{\mathrm{MLE}}=z=\left(z_{1},\ldots z_{p}\right).$
Then James - Stein Estimator is $\hat{\theta}^{\mathrm{JS}}=z-\frac{(p-2)z}{z^{\top}z}=z\left(1-\frac{p-2}{z^{\top}z}\right)=\hat{\theta}^{\mathrm{MLE}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{MLE}\top}\hat{\theta}^{\mathrm{MLE}}}\right).$
The risk of James - Stein Estimator equals to
$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$
$=\mathbb{E}\left(\left\{(z-\theta)-\frac{(p-2)z}{z^{\top}z}\right\}^{\top}\left\{(z-\theta)-\frac{(p-2)z}{z^{\top}z}\right\}\right)=$
$=\mathbb{E}\left((z-\theta)^{\top}(z-\theta)\right)-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}{\mathbb{E}\left(\frac{1}{z^{\top}z}\right)}=$
$=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)$
Let us consider the last two terms:
$-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
$=-2(p-2)\left(1-\mathbb{E}\left(\frac{z^{\top}\theta}{z^{\top}z}\right)\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
$=(p-2)\left(-2+\mathbb{E}\left(\frac{2z^{\top}\theta}{z^{\top}z}\right)+\mathbb{E}\left(\frac{(p-2)}{z^{\top}z}\right)\right)=(p-2)\left(-2+\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\right)$
We want to show that $\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)<2$
Let us denote $\tilde{x}=2z^{\top}\theta+p-2\quad$ and $\tilde{y}=\frac{1}{z^{\top}z}.$
By applying Cauchy–Schwarz inequality we get,
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)=\mathbb{E}(\tilde{x}\cdot\tilde{y})\leq|\mathbb{E}(\tilde{x}\cdot\tilde{y})|\leq\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}.$
Let us calculate each term separately:
$\mathbb{E}\left(\tilde{x}^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta+p-2\right)^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta\right)^{2}+4\left(z^{\top}\theta\right)(p-2)+(p-2)^{2}\right)=$
$=\mathbb{E}\left(4\left(x^{\top}\theta+y^{\top}\theta\right)^{2}+4\left(x^{\top}\theta+y^{\top}\theta\right)(p-2)\right)+(p-2)^{2}=$
$=4\mathbb{E}\left((x^{\top}\theta)^{2}\right)+8\mathbb{E}\left((x^{\top}\theta)(y^{\top}\theta)\right)+4\mathbb{E}\left((y^{\top}\theta)^{2}\right)+4(p-2)\mathbb{E}\left(x^{\top}\theta\right)+(p-2)^{2}=$
$=4Var\left(x^{\top}\theta\right)+4\left(\mathbb{E}\left(x^{\top}\theta\right)\right)^{2}+4Var\left(y^{\top}\theta\right)+4\left(\mathbb{E}\left(y^{\top}\theta\right)\right)^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=$
$=4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+4Var\left(y^{\top}\theta\right)+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}\leq$
$4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+8\epsilon^{2}||\theta||_{2}^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}.$
So, $\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\leq\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}.$
Note that,
$\mathbb{E}\left(z^{\top}z\right)=\mathbb{E}\left(\sum_{i}z_{i}^{2}\right)=\sum_{i}\mathbb{E}\left(z_{i}^{2}\right)=\sum_{i}Var(z_{i})+\sum_{i}\left(\mathbb{E}(z_{i})\right)^{2}\geq p+||\theta||_{2}^{2}.$
Using Jensen inequality, firstly for $f(x)=\frac{1}{x}$ and then for $g(x)=x^{2}$ we get
$\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}=\sqrt{\mathbb{E}\left(\frac{1}{(z^{\top}z)^{2}}\right)}\leq\sqrt{\frac{1}{\mathbb{E}\left((z^{\top}z)^{2}\right)}}\leq\frac{1}{\mathbb{E}\left(z^{\top}z\right)}\leq\frac{1}{p+||\theta||_{2}^{2}}.$
So, overall we get
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\leq\frac{\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}}{p+||\theta||_{2}^{2}}\leq 2.$
The final inequality can be established by direct computation. ∎
Theorem 3.3 Proof:.
[…]
Note that,
[…]
So, we need to show:
$c\left(2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right)+p\mathbb{E}\left(cV^{2}\right)\right)<0.$
We consider two cases depending on the sign of the constant $c$:
  • Case 1:
    $c<0$ and
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (7)
  • Case 2:
    $c>0$ and
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (8)
Let us derive some facts that will be helpful in the proof.
Firstly, we aim to show that for each $i$, the following inequality holds:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta.$
Note that:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)=1-\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}<0\right).$
Let us define:
$W_{i}:=\frac{2}{n}\sum_{j=1}^{n}(X_{ij}-\bar{X}_{i})(Y_{ij}-\bar{Y}_{i})=\frac{2}{n}\sum_{j=1}^{n}w_{j}.$
The centered random variables $X_{ij}-\bar{X}_{i}$ and $Y_{ij}-\bar{Y}_{i}$ are sub-Gaussian with parameters
$\sigma_{x,i}^{2}\left(1+\frac{1}{n}\right)\quad\text{and}\quad\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right),$
respectively. Their product $w_{j}$ is a sub-exponential random variable with parameters
$\nu^{2}=C_{1}\cdot\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2},\quad b=C_{2}\cdot\sqrt{\sigma_{x,i}^{2}\sigma_{y,i}^{2}\left(1+\frac{1}{n}\right)^{2}}.$
Reminder. A random variable $Z$ is sub-exponential if:
$\mathbb{E}[e^{\lambda Z}]\leq\exp\left(\frac{\lambda^{2}\nu^{2}}{2}\right),\quad|\lambda|<\frac{1}{b}.$
Bernstein Inequality for Sub-Exponential Random Variables. Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential variables with parameters $(\nu,b)$. Then for any $t>0$:
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\leq-t\right)\leq\begin{cases}\exp\left(-\dfrac{nt^{2}}{2\nu^{2}}\right),&\text{if }0\leq t\leq\dfrac{\nu^{2}}{b},\\ \exp\left(-\dfrac{nt}{2b}\right),&\text{if }t>\dfrac{\nu^{2}}{b}.\end{cases}$
Applying this inequality to $\frac{1}{n}\sum_{j}w_{j}$, we obtain:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}<0\right)=\mathbb{P}\left(\frac{2}{n}\sum_{j}w_{j}<-\hat{\sigma}^{2}_{y,i}\right)=\mathbb{P}\left(\frac{1}{n}\sum_{j}w_{j}<-\frac{\hat{\sigma}^{2}_{y,i}}{2}\right),$
which can be bounded by:
$\leq\begin{cases}\exp\left(-\dfrac{n(\hat{\sigma}^{2}_{y,i})^{2}}{8C_{1}\sigma_{x,i}^{2}\sigma_{y,i}^{2}(1+\frac{1}{n})^{2}}\right),&\text{if }\frac{\hat{\sigma}^{2}_{y,i}}{2}\leq\dfrac{\nu^{2}}{b},\\ \exp\left(-\dfrac{n\hat{\sigma}^{2}_{y,i}}{4C_{2}\sqrt{\sigma_{x,i}^{2}\sigma_{y,i}^{2}(1+\frac{1}{n})^{2}}}\right),&\text{otherwise}.\end{cases}$
Therefore, we conclude:
$\mathbb{P}\left(\hat{\sigma}^{2}_{y,i}+W_{i}\geq 0\right)\geq 1-\delta,$
where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$. Thus, the desired non-negativity condition holds with high probability in the large sample regime. We will use this fact in the proof of case 2. Now we are ready to consider each case separately.
Case 1. We examine the first term in more detail:
$2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]=2\sum_{i=1}^{p}\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}-\beta_{i}\right)V\right]$ (9)
$=2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)$
Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can apply Stein’s lemma for the Gamma distribution:
Lemma A.1 (Stein’s Lemma for the Gamma Distribution).
Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that
$\mathbb{E}[Xh(X)]<\infty\quad\text{and}\quad\mathbb{E}[Xh^{\prime}(X)]<\infty.$
Then the following identity holds:
$\mathbb{E}[(X-\alpha\beta)h(X)]=\beta\,\mathbb{E}[Xh^{\prime}(X)].$
We apply Stein’s identity to the function $h(x)=V$, where
$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}.$
Using the Gamma Stein identity:
$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)],$
we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ as follows:
$\frac{dV}{d\hat{\sigma}_{x,k}^{2}}=\frac{1}{p}\cdot V\cdot\frac{1}{\hat{\sigma}_{x,k}^{2}+\hat{\sigma}_{y,k}^{2}+w_{k}}.$
Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}}\right].$
In particular, we obtain the inequality:
$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$
Using this bound equation (9) becomes:
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}V\right]-\beta_{i}\cdot\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]+\mathbb{E}\left[\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right]\right)=$
$=2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$
Let’s denote
$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$
Overall, we get lower bound for :
$2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{0,i}-\beta_{i}\right)V\right]+pc\mathbb{E}[V^{2}]\geq-\frac{2p}{\alpha(\alpha+1)}\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]+pc\mathbb{E}[V^{2}]\geq$
$\geq-p\mathbb{E}\left[\overline{\hat{\sigma}_{z}^{2}}V\right]\left(\frac{2}{\alpha(\alpha+1)}+c\right)>0$
Last inequality appears due to assumption. So we have:
$c<0\ \text{and}\ \frac{2}{\alpha(\alpha+1)}+c<0\Rightarrow c\leq-\frac{2}{\alpha(\alpha+1)}.$
Case 2. We assume $c>0$ and we want to show
$02\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<$ 0 .
Recall that using Bernstein’s inequality we get with high probability, the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative. So, combining this with Stein’s lemma for $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ we get:
$\mathbb{E}(V)$ $=\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}\left(V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}\right)$
$\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}(V)$
$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$
$\Rightarrow\quad-\mathbb{E}(V)$ $\leq-\frac{1}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\cdot\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V).$
Plug in this bound and get inequality for  (9):
$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)$ $=2\sum_{i=1}^{p}\mathbb{E}\left(\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}-\beta_{i}\right)V+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}V\right)\right)\right)$
$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$
Using our previous notation $\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2},$
$=\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$
Overall for (8) we get:
$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})\leq$
$\leq\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)+pc\,\mathbb{E}(V^{2})<0$
The last inequality we want to establish by finding the appropriate $c$. So, we will rewrite it in a more suitable way:
$c\,\mathbb{E}(V^{2})<\frac{2}{\alpha p+1}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)-\frac{2}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)$
$c<\frac{2}{\alpha p+1}\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}-\frac{2}{\alpha+1}\cdot\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}$ (10)
Let’s bound each term independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, then we bound second term
$\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)\geq\mathbb{E}(V^{2})\quad\Rightarrow\quad-\frac{\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)}{\mathbb{E}(V^{2})}\leq-1.$ (11)
Also, using Cauchy–Schwarz inequality:
$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$
For the first term we get:
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}\cdot\sqrt{\mathbb{E}(V^{2})}}{\mathbb{E}(V^{2})}=\sqrt{\frac{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}{\mathbb{E}(V^{2})}}$ (12)
By applying Jensen’s inequality for $f(x)=\sqrt{x}$ – concave, we have:
$\sqrt{\mathbb{E}(V^{2})}\geq\mathbb{E}(|V|)=\mathbb{E}(V)\Rightarrow\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}$ (13)
Now,
$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq$
Again, using Jensen’s inequality for $f(x)=\exp{x}$
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$
We know that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
$\mathbb{E}(\log\sigma_{x,i}^{2})\ \geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$
Plug in this inequality to the derivations above we get:
$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$
For (13) we get:
$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$
Also, we can directly calculate numerator of (12):
Firstly,
$\mathbb{E}(\hat{\sigma}_{x,i}^{4})=Var(\hat{\sigma}_{x,i}^{2})+(\mathbb{E}(\hat{\sigma}_{x,i}^{2}))^{2}=\alpha\beta_{i}^{2}+\alpha^{2}\beta_{i}^{2}=\alpha\beta_{i}^{2}(1+\alpha).$
$\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2})=\alpha^{2}\beta_{i}\beta_{j}.$
$\sqrt{\mathbb{E}\left(\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\right)^{2}\right)}=\sqrt{\sum_{i=1}^{p}\mathbb{E}(\hat{\sigma}_{x,i}^{4})+2\sum_{1\leq i<j\leq p}\mathbb{E}(\hat{\sigma}_{x,i}^{2})\mathbb{E}(\hat{\sigma}_{x,j}^{2}})=$
$=\sqrt{\sum_{i=1}^{p}\alpha\beta_{i}^{2}(1+\alpha)+2\sum_{1\leq i<j\leq p}\alpha^{2}\beta_{i}\beta_{j}}=\sqrt{\alpha(1+\alpha)\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}\left((\sum_{i=1}^{p}\beta_{i})^{2}-\sum_{i=1}^{p}\beta_{i}^{2}\right)}=$
$=\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$
Therefor, for (12) we get upper bound:
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$
Using this inequality and (11) finally for (10) we get:
$c<\frac{2}{\alpha p+1}\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}-\frac{2}{\alpha+1}=$
$=\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$
Note, that
$\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}\geq\frac{p}{\sum_{i=1}^{p}\frac{1}{\beta_{i}}}\Rightarrow\frac{1}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\leq\frac{\sum_{i=1}^{p}\frac{1}{\beta_{i}}}{p}.$
Also, by Jensen’s inequality we get
$\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\beta_{i}}\leq\frac{1}{\frac{1}{p}\sum_{i=1}^{p}\beta_{i}}=\frac{p}{\sum_{i=1}^{p}\beta_{i}}.$
Hence,
$\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\leq\frac{1}{p}\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}\left(\sum_{i=1}^{p}\frac{1}{\beta_{i}}\right)$
$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$
So, dominance holds for
$0<c<\frac{2p}{(\alpha p+1)}\exp{\left(\frac{1}{\alpha}\right)}\sqrt{\frac{1}{\alpha}+1}-\frac{2}{\alpha+1}.$
Note that under regular conditions Stein’s estimate dominance holds when constant $c$ satisfies
$0<c<\frac{2p}{\alpha p+1}-\frac{2}{\alpha+1}=\frac{2(p-1)}{(\alpha+1)(\alpha p+1)}.$
In our current case, we instead obtain the condition
$0<c<\frac{2p}{\alpha p+1}\exp\left(\frac{1}{\alpha}\right)\sqrt{\frac{1}{\alpha}+1}-\frac{2}{\alpha+1},$
which yields a slightly higher upper bound.
Appendix B Experiments
We will, firstly, provide derivations of Lasso and Ridge shrinkage estimators for mean and variance.
Lasso mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda|\mu|\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{1}$ – penalty to the mean and the solution is given by the soft-thresholding operator:
$\hat{\mu}=sign(\bar{x})\max{(0,|\bar{x}|-\frac{\lambda}{2n}),}$
where $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ is the sample mean. This form encourages shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large.
Lasso variance shrinkage. A similar penalized formulation can be used for estimating the variance, leading to a soft-thresholding-type estimator for sample variance $s^{2}$:
$\hat{\sigma}^{2}=\max{(0,s^{2}-\frac{\lambda}{2}),}$
where $\lambda>0$ is regularization hyperparameter, $s^{2}=\frac{1}{n}\sum_{i}(x_{i}-\bar{x})^{2}.$
Ridge mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda\mu^{2}\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{2}$ – penalty to the mean. Taking the derivative and setting it to zero yields the following solution:
$\hat{\mu}=\frac{\sum_{i}x_{i}}{n+\lambda}.$
Ridge variance shrinkage. Similarly, we can obtain the ridge estimate for the variance:
$\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$
where $s^{2}$ is the sample variance and $\lambda>0$ is the regularization parameter.
Now let us discuss reproducibility of experiments on each dataset.
B.1 CIFAR10 dataset
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 4.
We used the standard ResNet-9 architecture, modifying only the batch normalization layers according to the method under consideration. For data splits, we used the standard predefined training and test sets. Training time was approximately $10$ minutes per run, with negligible variation across methods.
All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each experiment was repeated $11$ times to estimate the standard deviation of the performance metrics, and we tested three different batch sizes $(32,64,128)$, resulting in a total training time of approximately $1800$ minutes.
We used SGD with Nesterov acceleration and default parameters, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. We employed early stopping based on validation accuracy on data without noise to prevent overfitting. We used cross-entropy loss for training, and report classification accuracy as the main evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, we selected the regularization parameter $\lambda\in(10^{-3},10^{-1})$ yielding the best performance on validation dataset without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
Standard BN (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
Standard BN (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
Our BN (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our BN (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our BN (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Mean corrected BN (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Mean corrected BN (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
Mean corrected BN (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Khoshsirat et al. BN [15] (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat et al. BN [15] (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat et al. BN [15] (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
Lasso corrected BN (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
Lasso corrected BN (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
Lasso corrected BN (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
Ridge corrected BN(32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
Ridge corrected BN(64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
Ridge corrected BN(128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
Table 4: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
B.2 Cityscapes dataset
We performed segmentaion experiments on Cityscapes dataset – high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. In all our experiments we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 5. We use the HRNetV2-W16 architecture (where 16 denotes the width of the high-resolution convolution) as our baseline model. The original training set was randomly split into $80\%$ for training and $20\%$ for validation, while the original validation set was used as the test set. All images were resized from $1024\times 2048$ to $512\times 1024$ for computational efficiency. To account for variability and estimate the standard deviation of our results, each experiment was repeated $7$ times for each batch size $(16,32).$ Each run took approximately $1.5$ hours on a single NVIDIA A100 GPU with $64$ GB of RAM and $8$ CPU cores, resulting in a total training time of around 126 hours.
We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the mean Intersection over Union (mIoU) score on validation data without noise was employed to prevent overfitting. We used cross-entropy loss for training and report mIoU as the primary evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
Standard BN (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
Our BN (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our BN (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Mean corrected BN (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
Mean corrected BN (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
Khoshsirat et al. BN [15] (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsirat et al. BN [15] (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
Lasso corrected BN (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
Lasso corrected BN (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
Ridge corrected BN(8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
Ridge corrected BN(16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
Table 5: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
B.3 Details of Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN with Stein shrinkage on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). This data was obtained from the PPMI (Parkinson’s Progressive Markers Initiative) portal after a formal request was filled out. Please see data access citation and acknowledgment below for this data.
DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline involving Eddy currents and motion correction etc., using the FSL software [13]. DTIs are matrix-valued images in which each voxel takes on a value that is a $(3,3)$ symmetric positive definite (SPD) matrix representing the covariance matrix of a zero-mean local Gaussian at the voxel. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization is applied to these scalar feature maps to remove site effects caused by distinct acquisition protocols and equipment used. Harmonization was achieved using the method called ComBat adapted to diffusion tensor images and associated software presented in [12]. It should be noted that there are other sophisticated hamronization methods that one can apply to the data but this one is simple. The data set is imbalanced and consists of $490$ (130 PD and 360 Controls) FA maps. To address this, we use mixup-based[31] augmentation with sampling bias favoring the minority class (PD), this approach allowed us to generate additional training examples. We follow the same procedure in our other experiments for the setup and use a 5-fold cross validation to compare 3D CNN models with different variants of 3D BatchNorm layer. The architecture consists of four 3D convolutional layers, each followed by a 3D BatchNorm layer, and three linear layers for binary classification. The 3D input tensors have spatial dimensions of $(128\times 128)$ and a depth of $64$. We conduct experiments using two batch sizes ($10$ and $16$) and optimize the binary cross-entropy loss. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 6. We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the accuracy score on validation data without noise was employed to prevent overfitting. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise. All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each configuration take around $20$ minutes, which end up to the 20 hours for the whole experiments.
Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard BN (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our BN (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our BN (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Mean corrected BN (10) 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
Mean corrected BN (16) 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
Khoshsirat et al. BN [15] (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsirat et al. BN [15] (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso corrected BN (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso corrected BN (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge corrected BN(10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge corrected BN(16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
Table 6: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data. Values depict mean of accuracies $\pm$ standard deviation.